How do I describe how to solve the following using factorization and zero factor property.
1. x^2 -x - 6 = 0
Thank you!

Answer by algebrahouse.com(1079)

"How do I describe how to solve the following using factorization and zero factor property."

x^2 - x - 6 = 0
(x - 3)(x + 2) = 0 {factored into two binomials}
x - 3 = 0 or x + 2 = 0 {set each factor equal to 0}
x = 3 or x = -2 {solved each equation for x}
www.algebrahouse.com